Solve the differential equation 


dy/dx=x/16y

.

a) Find the implicit solution


b) Find the equation of the solution through the point (x,y)=(4,1) Your equation must describe a single curve of y=f(x) with the domain of f as large as possible. 


c) Find the equation of the solution through the point (x,y)=(0,−2) Your answer should be of the form y=f(x)



1
Expert's answer
2022-03-14T17:36:43-0400

a)..

"\text{To solve the separable equation } \frac{d y(x)}{d x}=\frac{1}{16} x y(x) :\\
\text{Divide both sides by } \frac{y(x)}{16} :\\"

"\frac{16 \frac{d y(x)}{d x}}{y(x)}=x\\[2mm]"

"\text{Integrate both sides with respect to }x :"


"\int \frac{16 \frac{d y(x)}{d x}}{y(x)} d x=\int x d x \\[4mm]"

"\text{Evaluate the integrals:}\\
16 \log (y(x))=\frac{x^{2}}{2}+c_{1}, \text{ where } c_{1} \text{ is an arbitrary constant.}"


"\text{Solve for } y(x):"


"y(x)=e^{1 / 32\left(x^{2}+2 c_{1}\right)}"

"\text{Simplifying the arbitrary constants, we have the implicit function to be:}\\

y(x)=e^{1 / 32\left(x^{2}+c_{1}\right)}"

b).


"\text{Solve for }c_{1} \text{using the initial conditions:}\\
\text{Substitute } y(4)=1 \text{ into } y(x)=e^{1 / 32\left(x^{2}+c_{1}\right)} : e^{1 / 32\left(c_{1}+16\right)}=1\\[4mm]
\text{Solving the equation:}\\
c_{1}=-16\\[4mm]

\text{Substitute } c_{1}=-16 \text{ into } y(x)=e^{1 / 32\left(x^{2}+c_{1}\right)} :\\[4mm]

\text{Answer:}\\

y(x)=e^{1 / 32\left(x^{2}-16\right)}"



c).

"\text{Solve for } c_{1} \text{ using the initial conditions:}\\
\text{Substitute } y(0)=-2 \text{ into } y(x)=e^{1 / 32\left(x^{2}+c_{1}\right)} : e^{c_{1} / 32}=-2\\[4mm]
\text{Solve the equation:}\\

c_{1}=32 i \pi+32 \log (2)\\[4mm]

\text{Substitute } c_{1}=32 i \pi+32 \log (2) \text{ into } y(x)=e^{1 / 32\left(x^{2}+c_{1}\right)} :\\[4mm]

\text{Answer:}\\
y(x)=-2 e^{x^{2} / 32}"